# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 4. Analysis of Cerami sequences
(4.17) $\begin{split}&\int_{\{u_{n}\geq 0\}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{\left(f(u_{n})\right)^{2}}\right)+(1-\tau)\int_{\{u_{n}<0\}}|\nabla u_{n}|^{2}\\ &\quad+\int_{\{u_{n}\geq 0\}}\left(\ln|\cdot|\ast F(u_{n})\right)F(u_{n})+(1-\tau)\int_{\{u_{n}<0\}}\left(\ln|\cdot|\ast F(u_{n})\right)f(u_{n})u_{n}\\ &=|\mathcal{I}^{\prime}(u_{n})[v_{n}]|\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|v_{n}\|\lesssim\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|u_{n}\|=\text{o}_{n}(1)\,.\end{split}$
Since $f\equiv 0$ on $\mathbb{R}^{-}$, the last term in the left-hand side is zero. Combining this with $\mathcal{I}(u_{n})\to c_{mp}$, one infers
$\begin{split}\|\nabla u_{n}\|_{2}^{2}-2c_{mp}+\text{o}_{n}(1)&=\int_{\mathbb{R}^{2}}\left(\ln\frac{1}{|\cdot|}\ast F(u_{n})\right)F(u_{n})\\ &=\int_{\{u_{n}\geq 0\}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{\left(f(u_{n})\right)^{2}}\right)+(1-\tau)\int_{\{u_{n}<0\}}|\nabla u_{n}|^{2}+\text{o}_{n}(1)\\ &\leq(1-\tau)\|\nabla u_{n}\|_{2}^{2}+\text{o}_{n}(1)\,,\end{split}$
from which it is easy to obtain
(4.18) $\begin{split}\|\nabla u_{n}\|_{2}^{2}\leq\frac{2c_{mp}}{\tau}+\text{o}_{n}(1)\,.\end{split}$
@@
(4.19) $|I_{0}(u_{n})|=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(|x-y|)F(u_{n}(x))F(u_{n}(y))\,{\rm d}x\,{\rm d}y\right|\leq C_{1}$
(4.20) $|A_{0}(F(u_{n}),f(u_{n})u_{n})|=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(|x-y|)F(u_{n}(x))f(u_{n}(y))u_{n}(y)\,{\rm d}x\,{\rm d}y\right|\leq C_{2}\,.$
Using (4.18), (4.19), and (4.20), the condition $\mathcal{J}(u_{n})\to 0$ directly implies
(4.21) $\begin{split}\|u_{n}\|_{p}^{2p}\leq C\left(\int_{\mathbb{R}^{2}}F(u_{n})\right)^{2}\leq C\,,\end{split}$
@@
Thanks to Lemma 4.3, from now on we can always suppose that Cerami sequences at level $c_{mp}$ verifying (4.14) are nonnegative. Indeed, $u_{n}^{-}:=\min\{u_{n},0\}\in X$ and $u_{n}^{-}\leq 0$ and thus, recalling that $f\equiv 0$ on $\mathbb{R}^{-}$ by assumption, one has
$\begin{split}\|\nabla u_{n}^{-}\|_{2}^{2}&=\|\nabla u_{n}^{-}\|_{2}^{2}+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u_{n})\right)f(u_{n})u_{n}^{-}\\ &=\mathcal{I}^{\prime}(u_{n})[u_{n}^{-}]\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|u_{n}^{-}\|=\text{o}_{n}(1)\end{split}$
since $\|u_{n}^{-}\|\leq\|u_{n}\|\leq C$ by Lemma 4.3. This implies that $u_{n}^{-}\to 0$ in $X$ as $n\to+\infty$ and therefore that $\{u_{n}^{+}\}_{n}$ is a Cerami sequence of $\mathcal{I}$ at level $c$, which henceforth we will simply denote by $\{u_{n}\}_{n}$.
@@ 4.1. Mountain pass level estimate in the critical case
Under assumption ($f_{sc}$) the boundedness of the Cerami sequences is enough to proceed with the main existence argument, see Section 5, since uniform estimates of the nonlinear terms may be deduced by (2.6) by choosing a suitably small exponent $\alpha$. This of course is not the case when we are dealing with critical exponential nonlinearities, and we need to prove that under ($f_{c}$), and in particular taking into account ($f_{4}$), the critical mountain pass level is below a noncompactness threshold. To this aim, let us introduce the usual Moser sequence
$\widetilde{w}_{n}(x)=\frac{1}{\sqrt{2\pi}}\begin{cases}\sqrt{\ln n}\ \ &\mbox{for}\ \,0\leq|x|\leq\frac{\rho}{n}\,,\\ \\ \displaystyle\frac{\ln(\rho/|x|)}{\sqrt{\ln n}}&\mbox{for}\ \,\frac{\rho}{n}<|x|<\rho\,,\\ \\ 0&\mbox{for}\ \,|x|\geq\rho\,.\end{cases}$
@@
$\begin{split}\|\widetilde{w}_{n}\|^{p}_{\ast,p}&=\int_{B_{\rho}}\ln(b+|x|)|\widetilde{w}_{n}|^{p}\,{\rm d}x\\ &=(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\int_{0}^{\rho/n}\ln(b+r)\,r\,{\rm d}r+\frac{(2\pi)^{1-\frac{p}{2}}}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,\ln(b+r)r\,{\rm d}r\\ &\leq(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\frac{\rho^{2}\ln(b+\rho/n)}{2n^{2}}+(2\pi)^{1-\frac{p}{2}}\frac{\ln(b+\rho)}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,r\,{\rm d}r\,.\end{split}$
@@
(4.22) $\delta_{n}=\rho^{\frac{4}{p}}\frac{(2\pi)^{\frac{2}{p}-1}}{\ln n}\ln^{\frac{2}{p}}(b+\rho)\left[\frac{[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)\right]^{\frac{2}{p}}+{\text{o}_{n}}\left(\frac{1}{\ln n}\right).$
@@
(4.24) $c_{mp}<\frac{2\pi}{\alpha_{0}}\,.$
@@
$\int_{\mathbb{R}^{2}}\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]F(t_{n}w_{n})\,{\rm d}x=\int_{B_{\rho}}\int_{|x-y|\leq 1}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))F(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\geq 0$
and thus a contradiction with (4.28).
@@
(4.29) $\begin{split}\int_{\mathbb{R}^{2}}&\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]t_{n}w_{n}f(t_{n}w_{n})\,{\rm d}x\\ &=\int_{\{|x|\leq\frac{\rho}{n},\,|y|\leq\frac{\rho}{n}\}}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))t_{n}w_{n}f(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\\ &\quad+\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\setminus\{|x|\leq\frac{\rho}{n},\,|y|\leq\frac{\rho}{n}\}}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))t_{n}w_{n}f(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\\ &=:T_{1}+T_{2}\,.\end{split}$
@@
By the very definition of $w_{n}$ (see (4.23)) and since $|x-y|<2\rho/n<1$, for $n$ large enough we can estimate $I_{1}$ as follows
$\begin{split}T_{1}&=\int_{B_{\rho/n}}t_{n}w_{n}f(t_{n}w_{n})\,{\rm d}y\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}F(t_{n}w_{n})\,{\rm d}x\\ &=\int_{B_{\rho/n}}t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\,f\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\,{\rm d}y\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,F\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\,{\rm d}x\\ &\geq 2\pi\beta\frac{{\rm e}^{\alpha_{0}t_{n}^{2}[\pi(1+\delta_{n})]^{-1}\ln n}}{2\alpha_{0}t_{n}^{2}[\pi(1+\delta_{n})]^{-1}\ln n}\int_{B_{\rho/n}}\!\!\!\,{\rm d}y\,\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,{\rm d}x\,.\end{split}$
@@
$\int_{B_{\rho/n}}\!\!\!\,{\rm d}y\,\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,{\rm d}x\geq|B_{\rho/n}|^{2}\ln\frac{n}{2\rho}=\pi^{2}\left(\frac{\rho}{n}\right)^{4}\ln\frac{n}{2\rho}$
@@
which is a contradiction, either if $t_{n}\to+\infty$ or $t_{n}$ stays bounded with $t_{n}^{2}\geq\frac{4\pi}{\alpha_{0}}+\delta_{0}$. The proof of Step 2 is then completed. Observe that, as a consequence of Step 1 and Step 2
@@
(4.32) $\frac{t_{n}^{2}}{1+\delta_{n}}\leq\frac{4\pi}{\alpha_{0}}+\frac{C}{\ln n}=\frac{4\pi}{\alpha_{0}}+{\rm{O}}\left(\frac{1}{\ln n}\right)\,.$
@@
$\begin{split}\left(\frac{4\pi}{\alpha_{0}}\right)^{2}+\text{o}_{n}(1)\geq t_{n}^{4}&\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{4\left(\frac{\alpha_{0}}{4\pi}\frac{t_{n}^{2}}{1+\delta_{n}}-1\right)\ln n}\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{-4\frac{\delta_{n}}{1+\delta_{n}}\ln n}\\ &\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{-\rho^{\frac{4}{p}}(2\pi)^{\frac{2}{p}-1}\ln^{\frac{2}{p}}(b+\rho)\left[\frac{[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)\right]^{\frac{2}{p}}+\text{o}_{n}(1)}.\end{split}$
@@
To avoid trivial solutions, in showing existence we will need to prove a result à la Lions (see Section 5 below). To this end, in the spirit of [14, Lemma 6.3], we need to improve the integrability for $F(u_{n})$, where $\{u_{n}\}_{n}$ is the Cerami sequence given by Lemma 4.2, since this will enable us to uniformly control the terms appearing from an application of the Hardy-Littlewood-Sobolev inequality. Here the mountain pass level estimate given by Lemma 4.5 plays a crucial rôle. Unlike [14, Lemma 6.3], we cannot apply Ruf’s version of the Trudinger-Moser inequality in $H^{1}(\mathbb{R}^{2})$ because of the lack of the mass term in (1.5); however, since our space $X\subset L^{p}(\mathbb{R}^{2})$, we will make use of the refinement of Cao’s inequality in Theorem 2.2.

### v2

@@ 4. Analysis of Cerami sequences
(4.17) $\int_{\{u_{n}\geq 0\}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{\left(f(u_{n})\right)^{2}}\right)+(1-\tau)\int_{\{u_{n}<0\}}|\nabla u_{n}|^{2}$
+$\quad+\int_{\{u_{n}\geq 0\}}\left(\ln|\cdot|\ast F(u_{n})\right)F(u_{n})+(1-\tau)\int_{\{u_{n}<0\}}\left(\ln|\cdot|\ast F(u_{n})\right)f(u_{n})u_{n}$
$=|\mathcal{I}^{\prime}(u_{n})[v_{n}]|\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|v_{n}\|\lesssim\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|u_{n}\|=\text{o}_{n}(1)$
by (4.14). Since $f\equiv 0$ on $\mathbb{R}^{-}$, the last term in the left-hand side is zero. Combining this with $\mathcal{I}(u_{n})\to c_{mp}$, one infers
$\|\nabla u_{n}\|_{2}^{2}-2c_{mp}+\text{o}_{n}(1)$ $=\int_{\mathbb{R}^{2}}\left(\ln\frac{1}{|\cdot|}\ast F(u_{n})\right)F(u_{n})$
=$=\int_{\{u_{n}\geq 0\}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{\left(f(u_{n})\right)^{2}}\right)+(1-\tau)\int_{\{u_{n}<0\}}|\nabla u_{n}|^{2}+\text{o}_{n}(1)$
$\leq(1-\tau)\|\nabla u_{n}\|_{2}^{2}+\text{o}_{n}(1)\,,$
from which
(4.18) $\|\nabla u_{n}\|_{2}^{2}\leq\frac{2c_{mp}}{\tau}+\text{o}_{n}(1)\,.$
@@
(4.19) $|I_{0}(u_{n})|\leq C_{1}\qquad\mbox{and}\qquad|A_{0}(F(u_{n}),f(u_{n})u_{n})|\leq C_{2}\,.$
Using (4.18) and (4.19), the condition $\mathcal{J}(u_{n})\to 0$ directly implies
(4.20) $\|u_{n}\|_{p}^{2p}\leq C\left(\int_{\mathbb{R}^{2}}F(u_{n})\right)^{2}\leq C\,,$
@@
Thanks to Lemma 4.3, from now on we can always suppose that Cerami sequences at level $c_{mp}$ verifying (4.14) are nonnegative. Indeed, $u_{n}^{-}:=\min\{u_{n},0\}\in X$ and $u_{n}^{-}\leq 0$, and thus, recalling that $f\equiv 0$ on $\mathbb{R}^{-}$ by assumption, one has
$\|\nabla u_{n}^{-}\|_{2}^{2}$ $=\|\nabla u_{n}^{-}\|_{2}^{2}+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u_{n})\right)f(u_{n})u_{n}^{-}$
$=\mathcal{I}^{\prime}(u_{n})[u_{n}^{-}]\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\|u_{n}^{-}\|=\text{o}_{n}(1)$
since $\|u_{n}^{-}\|\leq\|u_{n}\|\leq C$ by Lemma 4.3. This implies that $u_{n}^{-}\to 0$ in $X$ as $n\to+\infty$ and therefore that $\{u_{n}^{+}\}_{n}$ is a Cerami sequence of $\mathcal{I}$ at level $c_{mp}$, which henceforth we will simply denote by $\{u_{n}\}_{n}$.
@@ 4.1. Mountain pass level estimate in the critical case
Under assumption ($f_{sc}$), the boundedness of the Cerami sequences suffices to carry out the main existence argument, see Section 5, since uniform estimates of the nonlinear terms may be deduced by (2.7) by choosing a suitably small exponent $\alpha$. This of course is not the case when we are dealing with critical exponential nonlinearities, and we need to prove that under ($f_{c}$), and in particular taking into account ($f_{4}$), the critical mountain pass level is below a noncompactness threshold. To this aim, let us introduce the usual Moser sequence $\widetilde{w}_{n}:B_{\rho}\to\mathbb{R}^{+}$ defined by
$\widetilde{w}_{n}(x)=\frac{1}{\sqrt{2\pi}}\begin{cases}\sqrt{\ln n}\ \ &\mbox{for}\ \,0\leq|x|\leq\frac{\rho}{n}\,,\\ \\ \displaystyle\frac{\ln(\rho/|x|)}{\sqrt{\ln n}}&\mbox{for}\ \,\frac{\rho}{n}<|x|<\rho\,.\end{cases}$
@@
$\|\widetilde{w}_{n}\|^{p}_{\ast,p}$ $=(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\int_{0}^{\rho/n}\ln(b+r)\,r\,{\rm d}r+\frac{(2\pi)^{1-\frac{p}{2}}}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,\ln(b+r)\,r\,{\rm d}r$
$\leq(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\frac{\rho^{2}\ln(b+\rho/n)}{2n^{2}}+(2\pi)^{1-\frac{p}{2}}\frac{\ln(b+\rho)}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,r\,{\rm d}r\,.$
@@
(4.21) $\delta_{n}:=\rho^{\frac{4}{p}}\frac{(2\pi)^{\frac{2}{p}-1}}{\ln n}\ln^{\frac{2}{p}}(b+\rho)\left[\frac{[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)\right]^{\frac{2}{p}}.$
@@
(4.23) $c_{mp}<\frac{2\pi}{\alpha_{0}},$
where $\alpha_{0}$ is defined in $(f_{c})$.
@@
$\int_{\mathbb{R}^{2}}\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]F(t_{n}w_{n})\,{\rm d}x=\int_{B_{\rho}}\int_{|x-y|\leq 1}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))F(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\geq 0\,,$
a contradiction with (4.27).
@@
(4.28) $\int_{\mathbb{R}^{2}}$ $\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]t_{n}w_{n}f(t_{n}w_{n})\,{\rm d}x$
=$=\int_{\{|x|\leq\frac{\rho}{n},\,|y|\leq\frac{\rho}{n}\}}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))t_{n}w_{n}f(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\setminus\{|x|\leq\frac{\rho}{n},\,|y|\leq\frac{\rho}{n}\}}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))t_{n}w_{n}f(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y$
$=:T_{1}+T_{2}\,.$
@@
By the definition of $w_{n}$ (see (4.22)) and since $|x-y|<2\rho/n<1$, for $n$ large enough we can estimate $T_{1}$ as follows
$T_{1}$ $=\int_{B_{\rho/n}}t_{n}w_{n}f(t_{n}w_{n})\left(\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}F(t_{n}w_{n})\,{\rm d}x\right)\,{\rm d}y$
$=\int_{B_{\rho/n}}t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\,f\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\left(\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,F\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\,{\rm d}x\right)\,{\rm d}y$
$\geq 2\pi\beta\frac{{\rm e}^{\alpha_{0}t_{n}^{2}[\pi(1+\delta_{n})]^{-1}\ln n}}{2\alpha_{0}t_{n}^{2}[\pi(1+\delta_{n})]^{-1}\ln n}\int_{B_{\rho/n}}\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,{\rm d}x\,{\rm d}y\,.$
@@
$\int_{B_{\rho/n}}\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,{\rm d}x\,{\rm d}y\geq|B_{\rho/n}|^{2}\ln\frac{n}{2\rho}=\pi^{2}\left(\frac{\rho}{n}\right)^{4}\ln\frac{n}{2\rho}\,,$
@@
which is a contradiction, either if $t_{n}\to+\infty$ or $t_{n}$ stays bounded with $t_{n}^{2}\geq\frac{4\pi}{\alpha_{0}}+\delta_{0}$. The proof of Step 2 is then complete. Observe that, as a consequence of Step 1 and Step 2
@@
$\frac{t_{n}^{2}}{1+\delta_{n}}\leq\frac{4\pi}{\alpha_{0}}+\frac{C}{\ln n}=\frac{4\pi}{\alpha_{0}}+{\rm{O}}\left(\frac{1}{\ln n}\right)\,.$
@@
$\left(\frac{4\pi}{\alpha_{0}}\right)^{2}+\text{o}_{n}(1)\geq t_{n}^{4}$ $\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{4\left(\frac{\alpha_{0}}{4\pi}\frac{t_{n}^{2}}{1+\delta_{n}}-1\right)\ln n}\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{-4\frac{\delta_{n}}{1+\delta_{n}}\ln n}$
$\geq\frac{\pi^{3}\rho^{4}\beta}{\alpha_{0}}\,{\rm e}^{-\rho^{\frac{4}{p}}(2\pi)^{\frac{2}{p}-1}\ln^{\frac{2}{p}}(b+\rho)\left[\frac{[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)\right]^{\frac{2}{p}}+\text{o}_{n}(1)}.$
@@
To avoid trivial solutions, in showing existence we will need to prove a result à la Lions (see Section 5 below). To this end, in the spirit of [14, Lemma 6.3], we need to improve the integrability for $F(u_{n})$, where $\{u_{n}\}_{n}$ is the Cerami sequence given by Lemma 4.2, since this will enable us to uniformly control the terms appearing from an application of the Hardy-Littlewood-Sobolev inequality. Here the mountain pass level estimate given by Lemma 4.5 plays a crucial rôle. Unlike [14, Lemma 6.3], we cannot apply Ruf’s version of the Trudinger-Moser inequality in $H^{1}(\mathbb{R}^{2})$ because of the lack of the mass term in (1.5); however, since our space $X\subset L^{p}(\mathbb{R}^{2})$ we will make use of the refinement of Cao’s inequality in Theorem 2.2.
Lemma 4.6.
Assume $(f_{1})-(f_{4})$ and $(f_{c})$. Let $\{u_{n}\}_{n}\subset X$ be a nonnegative Cerami sequence for $\mathcal{I}$ at level $c_{mp}<\frac{2\pi}{\alpha_{0}}$, which is bounded in $X$. Then, the sequence
$v_{n}:=G(u_{n})\,,\qquad\mbox{where}\quad G(t):=\int_{0}^{t}\sqrt{\frac{F(s)f^{\prime}(s)}{f(s)^{2}}}\,{\rm d}s\,,$
has the following properties:
  1. i)
    $\|v_{n}\|_{p}\leq\|u_{n}\|_{p}$ and $\|v_{n}\|_{*,p}\leq\|u_{n}\|_{*,p}$;
  2. ii)
    $\{v_{n}\}_{n}\subset X$ and
    (4.33) $\|\nabla v_{n}\|_{2}^{2}=2c_{mp}+\text{o}_{n}(1)<\frac{4\pi}{\alpha_{0}}\,.$
  3. iii)
    $v_{n}>\sqrt{\tau}\,u_{n}$ a.e. in $\mathbb{R}^{2}$;
  4. iv)
    for all $\varepsilon>0$ there exists $t_{\varepsilon}>0$ such that
    (4.34) $u_{n}\leq t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\quad\ \mbox{for a.e.}\ \ x\in\mathbb{R}^{2}.$
Proof.
Note that $G$ is well-defined and $C^{1}$ thanks to ($f_{2}$). By ($f_{2}$) one has
and, by $(f_{1})$ and Hölder’s inequality,
$|G(t)|^{p}$ $\leq t^{\tfrac{p}{2}}\left(\int_{0}^{t}\frac{F(s)f^{\prime}(s)}{f(s)^{2}}\,{\rm d}s\right)^{\frac{p}{2}}\leq t^{\tfrac{p}{2}}\left(-\frac{F(t)}{f(t)}+\lim_{s\to 0^{+}}\frac{F(s)}{f(s)}+t\right)^{\frac{p}{2}}$
$\leq t^{p}\left(-\frac{F(t)}{tf(t)}+1\right)^{\frac{p}{2}}\leq t^{p}.$
This yields at once $\|v_{n}\|_{p}\leq\|u_{n}\|_{p}$ and $\|v_{n}\|_{*,p}\leq\|u_{n}\|_{*,p}$. Note also that $G(t)>\sqrt{\tau}t$, hence $v_{n}>\sqrt{\tau}\,u_{n}$ a.e. in $\mathbb{R}^{2}$. Recalling now that $u_{n}\geq 0$, and combining (4.17) with $\mathcal{I}(u_{n})\to c_{mp}$, we infer
$2c_{mp}+\text{o}_{n}(1)$ $=\|\nabla u_{n}\|_{2}^{2}+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u_{n})\right)F(u_{n})$
$=\|\nabla u_{n}\|_{2}^{2}-\int_{\mathbb{R}^{2}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{f(u_{n})^{2}}\right)=\|\nabla v_{n}\|_{2}^{2}\,.$
Since $c_{mp}<\frac{2\pi}{\alpha_{0}}$ by assumption, we deduce (4.33). Finally, using ($f_{3}$) and arguing as in [14, Lemma 6.3], one proves also (4.34). ∎